# Nonsingular black holes as dark matter

### v1

@@ Abstract
Introduction–The story of dark matter began in 1933 with Fritz Zwicky’s observation that the visible matter in the Coma cluster of galaxies was insufficient to account for its observed gravitational effects Zwicky (1933). In the 1970s, Vera Rubin’s studies of galactic rotation curves provided further compelling evidence Rubin et al. (1980). She discovered that stars in the outer regions of galaxies were orbiting at unexpectedly high speeds, challenging Newtonian predictions and suggesting the presence of unseen mass surrounding the galaxies. Additional support for dark matter comes from gravitational lensing, galaxy rotation curves, the cosmic microwave background and the formation of cosmic structures. Assuming general relativity, dark, non-baryonic matter is estimated to comprise nearly 25% of the total matter content of the universe Abbott et al. (2023). Despite decades of research and significant investment in direct detection experiments—including searches for weakly interacting massive particles (WIMPs) and axions—no conclusive evidence of dark matter particles has yet been found.
Many studies have shown that primordial black holes (PHBs) can be excellent dark matter candidates (see, e.g. Green and Kavanagh (2021); Villanueva-Domingo et al. (2021); Carr and Green (2024)). Stephen Hawking’s discovery that black holes radiate via quantum processes near the event horizon introduced the idea that black holes are not perfectly “black” but instead emit radiation by losing mass, leading to several important questions about what happens as they shrink in size. As the mass decreases, the temperature increases, leading to even higher-intensity radiation and eventually reaching a point where the black hole should evaporate entirely, leaving nothing behind. This is predicted to result in a thermal spectrum of particles emitted as the black hole disappears Hawking (1974).
Conventional calculations for a black hole of mass $m$ lead to a Hawking temperature, $T_{H}\propto 1/m$, with luminosity, $P\propto 1/m^{2}$. A solar mass black hole takes more than $10^{67}$ years to evaporate–significantly longer than the current age of the universe ($1.38\times 10^{10}$ years), while for a black hole of $10^{11}$ kg, the evaporation time is $2.6\times 10^{9}$ years. Thus, experiments are searching for signs of exploding primordial black holes. No such radiation has been observed. The evaporation observations place tight constraints on the viability of low-mass PBHs as dark matter, see yellow region of Fig. 1.
@@
In this paper, we demonstrate the infinity in the Hawking temperature is inextricably linked to another famous infinity: the singularity at the center of the black hole interior. We show how nonsingular black holes can naturally evolve to form stable remnants after the evaporation process. Indeed, early proposals that black hole evaporation resulted in Planck mass remnants argued that remnants were necessary for unitary time evolution, and hypothesized that these remnants could be dark matter Aharonov et al. (1987); MacGibbon (1987). Many others followed Barrow (1992); Barrow et al. (1992); Adler et al. (2001); Chen and Adler (2003); Carr and Kuhnel (2022); Profumo (2024).
@@
2D dilaton gravity toy model–To explore our hypothesis, consider a simplified and tractable 2D model. In the context of $(1+1)$ dilaton gravity, one may construct a nonsingular Schwarzschild-de Sitter (SdS) black hole metric Easson (2018) by utilizing an extremal curvature conjecture, which combines the maximal curvature conjecture of Markov (1982) with the minimal curvature conjecture of Easson (2007). The metric for this dilaton gravity black hole (DGBH) may be written:
@@
where $m$ is the mass of the black hole, $l$ is the fundamental length scale, and $\Lambda$ is the cosmological constant. This spacetime describes a singularity-free black hole in a de Sitter background. The solution has two horizons: a black hole horizon and a cosmological horizon. An observer passing through the black hole horizon approaches a smooth, constant, maximally curved spacetime, with Ricci scalar, $R_{max}\propto l^{-2}$. Outside of this horizon the solution resembles the familiar Schwarzschild-de Sitter solution and particle detectors respond to the thermal effects of both horizons.
Hawking radiation of nonsingular black holes–For our current investigation we consider Eq. 2, with $\Lambda=0$. To study the Hawking radiation process, we use the trace anomaly to obtain the flux of energy due to Hawking radiation Christensen and Fulling (1977). The trace of the anomalous energy-momentum tensor $T^{\mu}_{\mu}$ may be written in terms of curvature invariants and several cases in 2D are analyzed in Drehmer et al. (2009). In $(1+1)$ dimensions, the anomaly is expressed solely in terms of the Ricci scalar $R$, the only independent curvature invariant in 2D Deser et al. (1976):
@@
As expected, the remnant mass is proportional to the minimum length scale, $l$. Conventional MKS units are restored by multiplying the mass by $\nicefrac{{c^{2}}}{{G}}$. Setting $l$ to the Planck length, $l_{p}=1.6\times 10^{-35}$ $m$, yields,
@@
4D nonsingular black holes–Having established our proof of concept, we now move to four dimenensional models. We consider spherically symmetric metrics:
@@
The spacetimes are asymptotically flat and $m$ is the ADM mass. The Hawking temperatures may be calculated in terms of the surface gravity $\kappa$, as illustrated in the 2D case above. We observe that the qualitative features of nonsingular models exhibit a common trend: they reach a maximum temperature, which gradually approaches zero, likely leaving behind a remnant mass as in the the 2D DGBH case (see Fig. 2).
@@
A new window for PBH dark matter–We propose that dark matter may consist of nonsingular black holes and remnants formed in the primordial universe. These PBHs could have originated after inflation, during radiation domination, and before Big-Bang nucleosynthesis, from the collapse of large density perturbations and may constitute a significant fraction of the current matter-energy density if they formed as massive objects on the order of solar to multi-solar masses. Alternatively, if smaller, asteroid-sized PBHs formed in the range of $10^{17}\,\text{kg}$ to $10^{21}\,\text{kg}$, they could account for all of the observed dark matter Green (2024). A reasonable range for our fundamental length scale extends from the Planck scale to distances on the order of $10^{-19}\,\text{m}$, the latter corresponding to the TeV energy scale currently probed by the Large Hadron Collider (LHC). Consequently, for our black hole remnants, $10^{-35}\,\text{m}\leq l\leq 10^{-19}\,\text{m}$.
Primordial black holes can form during the early universe from the collapse of overdense regions. Such fluctuations generally have the form:
@@
Our nonsingular black holes do not evaporate fully. This significantly extends the lower mass limit, allowing $m_{\text{lower}}\approx l$. By choosing $m_{\text{lower}}\approx m_{\text{Planck}}$, corresponding to $l\approx 10^{-35}\,\text{m}$, and $m_{\text{upper}}\approx 10^{22}\,\text{kg}$, we find $\Omega_{\text{PBH}}^{\text{today}}\approx 0.23$, with the parameters from Eq. (22) falling within the expected range Carr (1975). Therefore, near Planck-mass nonsingular black holes or remnants could constitute all of the observed dark matter today, evading observational constraints set by conventional radiation calculations.
In addition to being dark matter candidates, nonsingular black holes have the potential to address other deep mysteries, such as the information loss paradox Polchinski (2017). They may provide baby universe interiors, or some other place to store in-falling information Easson and Brandenberger (2001). This work should be viewed as a foundational and admittedly speculative exploration. The results are primarily theoretical but potentially profound. The stability and observational characteristics of these remnants depend on the existence of a minimal length scale, and the effective field theories that prevent singularity formation, which likely violate traditional energy conditions Hawking and Penrose (1970); Zaslavskii (2010); Maeda (2022). Future advances in quantum gravity and observational techniques will be essential to test the robustness of nonsingular black holes as dark matter candidates.
@@ Acknowledgements.
We thank M. Baumgart, B. Carr, M. Parikh and B. Ratra for valuable discussions and correspondence. DAE is supported in part by the U.S. Department of Energy, Office of High Energy Physics, under Award Number DE-SC0019470.

### v2

@@ Abstract
I Introduction
The story of dark matter began in 1933 with Fritz Zwicky’s observation that the visible matter in the Coma cluster of galaxies was insufficient to account for its observed gravitational effects Zwicky (1933). In the 1970s, Vera Rubin’s studies of galactic rotation curves provided further compelling evidence Rubin et al. (1980). She discovered that stars in the outer regions of galaxies were orbiting at unexpectedly high speeds, challenging Newtonian predictions and suggesting the presence of unseen mass surrounding the galaxies. Additional support for dark matter comes from gravitational lensing, galaxy rotation curves, the cosmic microwave background and the formation of cosmic structures. Assuming general relativity, dark, non-baryonic matter is estimated to comprise nearly 25% of the total matter content of the universe Abbott et al. (2023). Despite decades of research and significant investment in direct detection experiments—including searches for weakly interacting massive particles (WIMPs) and axions—no conclusive evidence of dark matter particles has yet been found.
Many studies have shown that primordial black holes (PBHs) can be excellent dark matter candidates (see, e.g. Belotsky et al. (2014); Green and Kavanagh (2021); Villanueva-Domingo et al. (2021); Carr and Green (2024)). Stephen Hawking’s discovery that black holes radiate via quantum processes near the event horizon introduced the idea that black holes are not perfectly “black” but instead emit radiation by losing mass, leading to several important questions about what happens as they shrink in size. As the mass decreases, the temperature increases, leading to even higher-intensity radiation and eventually reaching a point where the black hole should evaporate entirely, leaving nothing behind. This is predicted to result in a culminating burst of high-energy particles as the black hole disappears. Hawking (1974).
Conventional calculations for a black hole of mass $m$ lead to a Hawking temperature, $T_{H}\propto 1/m$, with luminosity, $P\propto 1/m^{2}$. A solar mass black hole takes more than $10^{67}$ years to evaporate–significantly longer than the current age of the universe ($1.38\times 10^{10}$ years), while for a black hole of $10^{11}$ kg, the evaporation time is $2.6\times 10^{9}$ years. Thus, experiments are searching for signs of exploding primordial black holes. To date, no such radiation has been observed. The evaporation observations place tight constraints on the viability of low-mass PBHs as dark matter, see yellow region of Fig. 1.
@@ I Introduction
In this paper, we demonstrate that the infinity in the Hawking temperature is inextricably linked to another famous infinity: the singularity at the center of the black hole interior. We show how nonsingular black holes can naturally evolve to form stable remnants after the evaporation process. Indeed, early proposals that black hole evaporation resulted in Planck mass remnants argued that remnants were necessary for unitary time evolution, and hypothesized that these remnants could be dark matter Aharonov et al. (1987); MacGibbon (1987). Many others followed Barrow (1992); Barrow et al. (1992); Adler et al. (2001); Easson (2003); Chen and Adler (2003); Cai and Easson (2010); Dymnikova and Khlopov (2015); Carr and Kuhnel (2022); Profumo (2024); Calzà et al. (2025a, b).
@@
II 2D dilaton gravity model
To explore our hypothesis, consider a simplified and tractable 2D model. In the context of $(1+1)$ dilaton gravity, one may construct a nonsingular Schwarzschild-de Sitter (SdS) black hole metric Easson (2018) by utilizing an extremal curvature conjecture, which combines the maximal curvature conjecture of Markov (1982) with the minimal curvature conjecture of Easson (2007). The metric for this dilaton gravity black hole (DGBH) may be written:
@@
where $m$ is the mass of the black hole, $l$ is the fundamental length scale, and $\Lambda$ is the cosmological constant. This spacetime describes a singularity-free black hole in a de Sitter background. The solution has two horizons: a black hole horizon and a cosmological horizon. An observer falling through the black hole horizon approaches a smooth, constant, maximally curved spacetime, with Ricci scalar, $R_{max}\propto l^{-2}$. Outside of this horizon the solution resembles the familiar Schwarzschild-de Sitter solution and particle detectors respond to the thermal effects of both horizons.
III Hawking radiation of nonsingular black holes
For our current investigation we consider Eq. 2, with $\Lambda=0$. To study the Hawking radiation process, we use the trace anomaly to obtain the flux of energy due to Hawking radiation Christensen and Fulling (1977). The trace of the anomalous energy-momentum tensor $T^{\mu}_{\mu}$ may be written in terms of curvature invariants and several cases in 2D are analyzed in Drehmer et al. (2009). In $(1+1)$ dimensions, the anomaly is expressed solely in terms of the Ricci scalar $R$, the only independent curvature invariant in 2D Deser et al. (1976):
@@ III Hawking radiation of nonsingular black holes
As expected, the remnant mass is proportional to the minimum length scale, $l$. Conventional MKS units are restored by multiplying the mass by $\frac{c^{2}}{G}$. Setting $l$ to the Planck length, $l_{p}=1.6\times 10^{-35}$ $m$, yields,
@@
IV 4D nonsingular black holes
Having established our proof of concept, we now move to four dimensional models. We consider spherically symmetric metrics:
@@
The spacetimes are asymptotically flat and $m$ is the ADM mass. The Hawking temperatures may be calculated in terms of the surface gravity $\kappa$, as illustrated in the 2D case above. We observe that the qualitative features of nonsingular models exhibit a common trend: the black hole temperature reaches a maximum and then gradually falls to zero, likely leaving behind a remnant mass as in the 2D DGBH case (see Fig. 2). [^1] Yet another way such remnants can form in quantum gravity arises in the context of asymptotically safe gravity Cai and Easson (2010)..
@@
V New window for PBH dark matter
We propose that dark matter may consist of nonsingular black holes and their asymptotic remnants formed in the primordial universe. These PBHs could have originated after inflation, during radiation domination, and before Big-Bang nucleosynthesis, from the collapse of large density perturbations and may constitute a significant fraction of the current matter-energy density if they formed as massive objects on the order of solar to multi-solar masses. Alternatively, if smaller, asteroid-sized PBHs formed in the range of $10^{17}\,\text{kg}$ to $10^{21}\,\text{kg}$, they could account for all of the observed dark matter Green (2024). A reasonable range for our fundamental length scale extends from the Planck scale to distances on the order of $10^{-19}\,\text{m}$, the latter corresponding to the TeV energy scale currently probed by the Large Hadron Collider (LHC). Consequently, for our black hole remnants, $10^{-35}\,\text{m}\leq l\leq 10^{-19}\,\text{m}$.
V.1 Dark Matter Abundance of Remnants
Next, we estimate whether these remnants could constitute the dark matter of the universe. Primordial black holes can form during the early universe from the collapse of overdense regions. Such density fluctuations generally have the form:
@@ V.1 Dark Matter Abundance of Remnants
Our nonsingular black holes do not fully evaporate. This significantly extends the lower mass limit, allowing $m_{\text{lower}}\approx l$. By choosing $m_{\text{lower}}\approx m_{\text{Planck}}$, corresponding to $l\approx 10^{-35}\,\text{m}$, and $m_{\text{upper}}\approx 10^{22}\,\text{kg}$, we find $\Omega_{\text{PBH}}^{\text{today}}\approx 0.23$, with the parameters from Eq. (22) falling within the expected range Carr (1975). Therefore, near Planck-mass nonsingular black holes or remnants could constitute all of the observed dark matter today, evading observational constraints set by conventional radiation calculations.
VI Conclusions
In addition to being dark matter candidates, nonsingular black holes have the potential to address other deep mysteries, such as the information loss paradox Polchinski (2017). They may provide baby universe interiors, or some other place to store in-falling information Easson and Brandenberger (2001). This work should be viewed as a foundational and admittedly speculative exploration. The results are primarily theoretical but potentially profound. The stability and observational characteristics of these remnants depend on the existence of a minimal length scale, and the effective field theories that prevent singularity formation, which likely violate traditional energy conditions Hawking and Penrose (1970); Zaslavskii (2010); Maeda (2022) (see Appendix A). Future advances in quantum gravity and observational techniques will be essential to test the robustness of nonsingular black holes as dark matter candidates.
@@ Acknowledgements.
We thank M. Baumgart, B. Carr, M. Parikh, B. Ratra and S. Vagnozzi for valuable discussions and correspondence. DAE is supported in part by the U.S. Department of Energy, Office of High Energy Physics, under Award Number DE-SC0019470.
Appendix A Properties of supporting matter
The Birkhoff theorem establishes that the singular Schwarzschild black hole metric is the unique spherically symmetric vacuum solution to the Einstein field equations Birkhoff (1923)—a profound result that also extends to Kleinian (split-signature) metrics Easson and Pezzelle (2024). The theorem ensures that any nonsingular spherically symmetric metric, including those studied here, must be supported by matter in standard Einstein gravity.
Natural questions arise such as, what is the origin of this supporting matter which ultimately must violate traditional energy conditions to prevent singularity formation? How does this matter interact with ordinary matter from the Standard Model? Ultimately the interactions of this unusual supporting matter must be strongly constrained to avoid giving rise to long range ‘fifth forces’ and violations of the equivalence principle. The theories must obey solar system tests of gravity. While the properties of this supporting matter are exotic, at the classical level the alternative to this matter is formation of a singularity in gravitational collapse, which is arguably the most pathological option in a physical theory. In this appendix, we explore the nature of this supporting matter, beginning with an analysis of the energy conditions it must satisfy or violate.  [^2] We do not analyze the energy conditions for the metric Eq. 2, as the model is formulated in 2D dilaton gravity, where known subtleties arise in defining the energy conditions Chatterjee et al. (2013)..
A.1 Energy Conditions
It is well known that the traditional energy conditions must be violated by matter in order to avoid singularities Penrose (1965); Hawking (1966a, b, 1967); Hawking and Penrose (1970). The theorems identify the necessary criteria for establishing geodesic completeness in both stationary and time-dependent cosmological spacetimes Lesnefsky et al. (2023); Easson and Lesnefsky (2024a, b). We now examine the degree of the energy condition violation needed to support the nonsingular solutions of Table I.
The regular black holes are supported by effective energy momentum tensors of Type I S.W.Hawking and G.F.R.Ellis (1973), whose canonical form in the local Lorentz frame is
$T^{(\hat{a})(\hat{b})}=T^{\hat{\mu}\hat{\nu}}e^{(\hat{a})}_{\hat{\mu}}e^{(\hat{b})}_{\hat{\nu}}\,,$ (24)
where $\{e_{(\hat{a})}^{\hat{\mu}}\}$ with $a=0,1,2,3$, is a set of orthonormal basis vectors in the local Lorentz frame having $e_{(\hat{a})}^{\hat{\nu}}e_{(\hat{b})\hat{\nu}}=\eta_{(\hat{a})(\hat{b})}$. Here $\eta_{(\hat{a})(\hat{b})}$ is a Minkowski metric in the local Lorentz frame and the curved metric is $g_{\hat{\mu}\hat{\nu}}=\eta_{(\hat{a})(\hat{b})}e^{(\hat{a})}_{\hat{\mu}}e^{(\hat{b})}_{\hat{\nu}}$. Energy density and pressure are given by the compentns of the mixed Einstein tensor:
$T^{\hat{\mu}}{}_{\hat{\nu}}=G^{\hat{\mu}}{}_{\hat{\nu}}=diag(-\rho,p_{\parallel},p_{\perp},p_{\perp})\,,$ (25)
where we have taken the reduced Planck mass $M_{\rm Pl}^{-2}=8\pi G=1$. For the models considered here the energy density is $\rho=-T^{\hat{t}}{}_{\hat{t}}$ and we have pressures parallel to the radial coordinate, $p_{\parallel}=T^{\hat{r}}{}_{\hat{r}}$ and perpendicular to the radial coordinate, $p_{\perp}=T^{\hat{\theta}}{}_{\hat{\theta}}=T^{\hat{\phi}}{}_{\hat{\phi}}$. It is a further well-know property of our spacetimes that, $T^{\hat{r}}{}_{\hat{r}}=T^{\hat{t}}{}_{\hat{t}}$.
For the metric Eq. 19 with $M(r)$ in Eq. 20,
$\rho=-p_{\parallel}=\frac{2M^{\prime}(r)}{r^{2}}\,,\qquad p_{\perp}=-\frac{M^{\prime\prime}(r)}{r}\,.$ (26)
[…]
In most of our models, coordinates span the ranges $t\in(-\infty,\,\infty)$, $r\in(-\infty,\,\infty)$, $\theta\in[0,\,\pi]$ and $\phi\in(-\pi,\,\pi]$; For our purposes of studying energy conditions it is sufficient to focus our plots in the range $r\in[0,\,\infty)$; however, some of spacetimes such as Bardeen cover the full range $r\in(-\infty,\,\infty)$ (see, Fig. 9 of Easson et al. (2020).) For physically reasonable values of parameters resulting in positive mass and black hole horizons, we find qualitatively similar energy condition behaviors for Bardeen and Hayward spacetimes (see Fig. 3) and similarly for Dymnikova and Fan-Wang spacetimes (see Fig. 4).
[…]
From Fig. 3, (for the given parameters) we see the Bardeen and Hayward spacetimes obey the NEC and WEC for the entire domain of $r$, as depicted by the positivity of the solid red and green curves and the equality of the dotted red and yellow curves, $p_{\parallel}=-\rho$. The DEC is violated in the asymptotically flat region as seen from the blue curve superseding the red curve at large $r$. The SEC is violated (as expected) near the de Sitter core as seen by the negativity of the blue curve near $r=0$.
From Fig. 4, (for the given parameters) we see the Dymnikova and Fan-Wang spacetimes obey the NEC, WEC and DEC for the entire domain of $r$, as depicted by the positivity of the solid red and green curves and the equality of the dotted red and yellow curves, $p_{\parallel}=-\rho$. The DEC is satisfied as the blue curve remains bound by the red and dotted red curves for all $r$. Only the SEC is violated near the de Sitter core for $r\leq l$.
A.2 Matter Lagrangian
After evaluating the energy conditions that must be violated to sustain the regular black hole geometries in Table I, we now briefly discuss the type of matter Lagrangian required to source these spacetimes and how such spacetimes may form from gravitational collapse.
In Ayon-Beato and Garcia (1998, 2000) it was shown that the Bardeen regular black hole spacetime can be interpreted as an exact solution in general relativity with a source arising from nonlinear electrodynamics (NLED). A key characteristic of NLED models is that gauge field quantities remain finite everywhere, including at the location of a point source, inspiring the authors to explore the Bardeen black hole as a solution sourced by a monopole within an NLED framework.
The source is a magnetic monopole with charge $g$, with a Lagrangian given by
$\mathcal{L}=\frac{3m}{|g|^{3}}\Bigg{(}\frac{\sqrt{2g^{2}F}}{1+\sqrt{2g^{2}F}}\Bigg{)}^{5/2},$ (27)
where $F\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ and $m$ is the black hole mass. The ansatz for the gauge field and metric
$-g_{tt}(r)=1-\frac{2M(r)}{r},\ \ \ \ \ \ \ \ F_{\mu\nu}=2\delta^{\theta}_{[\mu}\delta^{\phi}_{\nu]}B(r,\theta),$ (28)
become solutions to the Einstein-NLED equations when
$M(r)=\frac{mr^{3}}{(r^{2}+g^{2})^{3/2}},\ \ \ \ \ \ \ B(\theta)=g\sin\theta.$ (29)
This form matches the Bardeen black hole with $g\rightarrow l$.
The other models of Table I may follow from similar NLED constructions. As there are several ways a black hole can be made regular using matter, the Birkoff theorem no longer applies and one may be concerned that the end state of general gravitational spherical collapse may no longer be unique. Might there be a physical reason to favor one regular black hole metric over another? It is intriguing that the Bardeen solution appears to emerge from fundamental principles in string theory such as T-duality Gaete et al. (2022).
A.3 Formation and observational signatures
Recently, in Bueno et al. (2024a, b) it is argued that dynamical gravitational collapse in a theory with an infinite tower of higher-derivative corrections to the Einstein-Hilbert action, under very general conditions, leads to the formation of regular black holes.
Finally, advances in the formation process of regular black holes has led to the exciting possibility of observational signatures of regular black holes, including features in regular black hole shadows from phase transitions during collapse potentially delaying apparent horizon formation Yan et al. (2020); Balart et al. (2025); Vertogradov et al. (2025).